OpenStudy (anonymous):

Could you please solve |4x-3| = 5 step by step, clearly showing and explaining as you go? Thank you in advance.

OpenStudy (anonymous):

|4x-3|=5 4x-3+3=5+3 4x=8 4x/4=8/4 x=2 There are two x's in this problem because of the || sign. That means absolute value. This is the other x. |4x-3|=5 4x+3=5 4x+3-3=5-3 4x=2 4x/4=2/4 x=1/2

OpenStudy (anonymous):

I went to http://www.myalgebra.com/ and entered your problem. It gave me x=2, -1/2

OpenStudy (anonymous):

I already know the solution. I just wanted to know how to get it.. and @abbyshorey, the answer you ended up doesn't match with solution that I'm already given?

OpenStudy (anonymous):

What was the solution they gave you?

OpenStudy (anonymous):

x = -1/2 (same as tagalong's)

OpenStudy (anonymous):

Oh ya! I forgot that absolute value problems has one negative in it. The work is still good though. Do you understand the steps?

OpenStudy (anonymous):

Ah, I see now. Thank you so much :)

OpenStudy (anonymous):

Anytime :)